A galvano meter gives a deflection of 55devesion when is shunt s of resistance of 8 ohm is connected in a parallel to it if it produces the deflection of 45 devesions find the resistance of galvano meter
1
Expert's answer
2021-08-06T15:26:05-0400

Given:

"I=55"

"I'=45"

"R_{\rm sh}=8\:\Omega"

Total current flowing through the both galvanometer and shunt is 

"I=I'+I_{\rm sh}"

For parallel connection of the wires

"I'R_A=I_{\rm sh}R_{\rm sh}"

Combining these equations give

"R_A=\frac{I-I'}{I'}R_{\rm sh}=\frac{55-45}{45}*8\:\Omega=1.78\:\Omega"